• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mercado, KaiPan

    Log entry time 16:16:08 on September 7,2009

    Entry number 288436

    keyword=Day Shift Summary

    SL: Luis Mercado (Kai Pan)
    TO: Huan Yao
    3rd: Buddhini Waidyawansa

    PS: Luis will be not available in the first a couple of hours during the shift, so Kai is helping to take the shift as a SL for Luis for the first a few hours.

    8:00 day shift start. Run # 12075, which is the last run in the owl shift, run end failed. Mark reboots the ROC23, and we start a new run #12076.

    9:00 Run # 12076 ended normally, however the online replay shows that the BPM12 saturated. End the run, and Richard is calling MCC to fix the problem.

    9:30 compton #18989 ended failed, reset and start a new run # 18990.

    9:30 BPM saturation problem fixed (see separate halog). start new run#
    12078 and online replay results looks good.


    11:21 Called MCC, BPM saturating. Fixed at event ~22000 of run 12080.

    11:25 BPM saturated again. Called MCC and asked them to lock RF Gain Setting to Low.

    11:41 BPM still stable after 15 minutes.

    11:58 BPM saturated again. For about 4 minutes while beam kept tripping. Asked them to lock RF Gain again.

    12:46 BPM still stable after 45 minutes!

    14:07 Still running smoothly...

    15:11 MCC having trouble delivering beam. Said it would be back in about 10-15 minutes.

    15:40 Beam is back but we are calling for a quick Controlled Access to do some Compton work.


    16:00 Shift over. Still in Controlled Acces.mercasdo